RIGOROUS RELATIVISTIC METHODS FOR ADDRESSING P- AND T -NONCONSERVATION IN HEAVY-ELEMENT MOLECULES

\underline{TIMO FLEIG}, Laboratoire de Chimie et Physique Quantiques, Université Paul Sabatier Toulouse 3, Toulouse, France.

A new and rigorous method for accurate ab-initio calculations of the electron electric dipole moment P, T -odd interaction constant is presented. The approach uses string-based Configuration Interaction wavefunctions^{a} and Dirac four-component spinors as one-particle basis functions, and the P, T -odd constant is obtained as an expectation value over these correlated wavefunctions. The method has been applied to the HfF^{+} molecular ion to determine spectroscopic constants for four low-lying electronic states. For one of these states (Ω = 1) we have determined a new accurate benchmark value^{b} for the effective electric field E_{eff} correlating 34 valence and outer atomic core electrons and using wavefunction expansions with nearly 5 · 10^{8} coefficients. For the Ω = 1 state of the ThO molecule the first ab-initio result for the electron EDM interaction constant is presented. Aspects of modern all-electron relativistic many-body approaches applicable to both atoms and molecules will be discussed, including perspectives for the treatment of other interesting candidate systems and P- or P, T -non-conserving effects in molecular systems.

a

S. Knecht, H. J. Å. Jensen and T. Fleig J. Chem. Phys. \underline{132}, 014108 (2010)

T. Fleig, H. J. Å. Jensen, J. Olsen and L. Visscher J. Chem. Phys. \underline{124}, 104106 (2006)

bT. Fleig and M. K. Nayak Phys. Rev. X \underline{XXX}, XXXX (submitted).